OpenStudy (anonymous):

--------------MEDAL AND FAN!!!!!---------- Find the sum or difference. (x^2 + 6x + 1) - (2x^2 - 8x + 4)

OpenStudy (usukidoll):

distribute the negative and combine like terms

OpenStudy (anonymous):

Umm how? What number do i distribute with?

OpenStudy (usukidoll):

right hand side of the equation distribute the negative sign

OpenStudy (anonymous):

So additive inverse? Switch the negative to addition and change the numbers to the right of it to its opposite?

OpenStudy (usukidoll):

\[ - (2x^2 - 8x + 4) = -2x^2+8x-4\]

OpenStudy (anonymous):

Oh ok. Whats the next step?

OpenStudy (usukidoll):

combine like terms and simplify

OpenStudy (usukidoll):

\[x^2+6x+1-2x^2+8x-4 = -x^2+14x-3\]

OpenStudy (anonymous):

Wait but i have a question. How do we combine like terms with exponents? For example: x^2 and 2x^2. How do we combine them?

OpenStudy (usukidoll):

-_______- by either subtracting or adding..

OpenStudy (anonymous):

Oh i get it. Sorry. Lol. Can you help me with another one?

OpenStudy (usukidoll):

eating

OpenStudy (anonymous):

Oh ok. Sorry. Thanks!